OpenStudy (blondey_315):

Solve for r: |5r-18|=0 3.6 (3.6, -3.6) No solution

OpenStudy (anonymous):

|5r-18|=0 5r = 18 r = 18/5 3.6

OpenStudy (anonymous):

just only 3.6

OpenStudy (blondey_315):

Thankss!

OpenStudy (anonymous):

you are welcome :)

OpenStudy (blondey_315):

I have like 7 questions left that I don't know how to do :(

OpenStudy (anonymous):

post on message :)

OpenStudy (anonymous):

The formula for the perimeter of a rectangle is P = 2l + 2w. If P = 104 and l = 13, what is the length of w? Answer: Put P = 104 and l = 13 into P = 2l + 2w P = 2l + 2w (104) = 2*(13) + 2W 104 = 26 + 2W 104 - 26 = 2W 78 = 2W 2W = 78 W = 78/2 W =39

OpenStudy (anonymous):

@blondey_315 here

OpenStudy (blondey_315):

ohh okayy

OpenStudy (anonymous):

what else?

OpenStudy (blondey_315):

Solve for b. (2b/4)+(b/3) 12 10 8 5

OpenStudy (anonymous):

\[\frac{ 2b }{ 4 } + \frac{ b }{ 3 } = \frac{ (3 \times 2b) + (4 \times b) }{ 12 } = \frac{ 6b + 4b }{ 12 } = \frac{ 10b }{ 12 }\]

OpenStudy (anonymous):

is in the questions (2b/4)+(b/3) = 0 ?

OpenStudy (anonymous):

or ?

OpenStudy (blondey_315):

is what in the question?

OpenStudy (anonymous):

would you show me the original text of that question?

OpenStudy (blondey_315):



OpenStudy (anonymous):

(2b/4)+(b/3) = X to find the value of b, we need to know what X is. We can't find b if it's only (2b/4)+(b/3)is known

OpenStudy (anonymous):

@blondey_315 are you still here?

OpenStudy (blondey_315):

yup

OpenStudy (anonymous):

hmm.., i don't know for this, let's go to the next question

OpenStudy (blondey_315):

wait it equals 10

OpenStudy (blondey_315):

(2b/4)+(b/3)=10, i think i put 0 last time i dont know would having 10 at the end help at all?

OpenStudy (anonymous):

@blondey_315

OpenStudy (blondey_315):

huh?

OpenStudy (blondey_315):

huh? @gerryliyana

OpenStudy (anonymous):

ok let we check

OpenStudy (anonymous):

\[\frac{ 2b }{ 4 } + \frac{ b }{ 3 } = 10\] \[\frac{ (3 \times 2b) + (4 \times b) }{ 12 } = 10\] \[\frac{ 6b + 4b }{ 12 } = 10\] \[\frac{ 10b }{ 12 } = 10\] \[10 b = 10 \times 12 = 120\] \[b = \frac{ 120 }{ 10 }\] \[b = 12\]

OpenStudy (anonymous):

@blondey_315 it's done

OpenStudy (blondey_315):

Yus that makes sense!

OpenStudy (blondey_315):

im sending it to you in DM so i can paste @gerryliyana

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

First 4n + 3 = 11 4n = 11-3 = 8 n = 8/4 = 2 Second 4(1 + n) = -4 4 + 4n = -4 4n = -4 -4 4n = -8 n = -8/4 n = - 2 Third 5 = 3n - 1 5 + 1 = 3n 6 = 3n n = 6/3 n = 2 4th 5(n + 1) = 2 5n + 5 = 2 5n = 2-5 5n = -3 n = -3/5

OpenStudy (anonymous):

the second equation has a solution of -2

OpenStudy (anonymous):

@blondey_315 get it?

OpenStudy (blondey_315):

yup i was looking it over it makes sense now thanks!

OpenStudy (blondey_315):

Solve |2y - 2| = 15 8.5 -13 8.5 and -6.5 17 and -13

OpenStudy (anonymous):

you are welcome :)

OpenStudy (blondey_315):

Yay it let me paste~

OpenStudy (anonymous):

|2y - 2| = 15 (|2y - 2|)^2 = 15^2 2y - 2 = + - sqrt(15^2) 2y - 2 = + - 15 So 2y - 2 = 15 2y = 15+2 = 17 y = 17/2 y = 8.5 or 2y - 2 = -15 2y = -15+2 2y = -13 y = -13/2 y = -6.5 So the answer are y = 8.5 and y = -6.5

OpenStudy (blondey_315):

ohhhhhh alright so if its in I I like that i square it?

OpenStudy (anonymous):

yes

OpenStudy (blondey_315):

ohhhhh i see!

OpenStudy (anonymous):

\[|2y - 2| = 15\] \[(|2y - 2|)^2 = 15^2\] \[2y - 2 = \pm 15\]

OpenStudy (anonymous):

you you will have 2 solutions

OpenStudy (blondey_315):

The formula for the volume of a rectangular box is V = LWH. If V = 200, L= 12.5, and W = 8, find the height of the box, to the nearest whole unit. 1 2 3 4

OpenStudy (anonymous):

using + or -

OpenStudy (blondey_315):

ohhh now i know and feel so stupid for not knowing before!

OpenStudy (blondey_315):

Thats the last question by the way!

OpenStudy (blondey_315):

I finished the test and got a 88 thanks! i wouldnt have gotten such a good grade with out your help!! :)

OpenStudy (anonymous):

V = LWH. If V = 200, L= 12.5, and W = 8 V = LWH. 200 = (12.5)*(8)*(H) H = 200/100 H = 2 the height of the box is 2.

OpenStudy (anonymous):

Congratulations!! :) Good job, @blondey_315 :)